# MindVote: When AI Meets the Wild West of Social Media Opinion

### v3

@@ 3.2 Experiment Design
4 Result
4.1 Performance of General Purpose Models
4.2 Survey-based Specialization Pitfalls
Our result in Table 2 reveals that fine-tuning on a structured survey dataset results in a significant performance drop on authentic, in-domain tasks. To demonstrate this, we compare our specialization models (base models that fine-tuned on traditional but context-stripped survey data) against their original base instruction-tuned models (the general-purpose models).
As shown in Table 2, the specialization models consistently underperform their base counterparts. This evaluation was conducted under our standard testing condition, where all models were provided with full context served as context priming setting. For instance, the SubPop-Llama-2-13B specialization model exhibits a 3.3 percentage point drop in its 1-Wass. score compared to the base Llama-2-13B model. This finding suggests that fine-tuning on sanitized data fails to equip models with the necessary capabilities to handle real-world contextual complexity. Instead of improving performance, this narrow specialization is detrimental, degrading the models’ ability to generalize to authentic social discourse.
@@ 4.3 Performance Across Topics and Cultures
Performance varies by Topic. As shown in Table 3, model performance is not uniform across different topical domains. On average, models achieve higher performance on Social Issues and Lifestyle topics, which often feature more structured discourse patterns that align with standard pre-training corpora. In contrast, performance tends to be lower in domains like Technology and Entertainment where specialized jargon, community-specific norms, and rapidly evolving vernacular are more common. This disparity suggests that models struggle to adapt their reasoning from general knowledge to the specific linguistic and social conventions of diverse online communities.
Performance Reflects Cultural Origin. The analysis in Figure 4 reveals a ’Cultural Gap,’ showing models have a strong cultural alignment with their origin. Western models like Gemini-2.5-Pro excel on Reddit, while Chinese models such as DeepSeek-R1 dominate on Weibo. We confirmed this gap is primarily cultural, not linguistic, by evaluating translated content; the performance penalty from translation was minimal compared to the large drop across cultural sources. This systematic variation demonstrates a clear alignment effect, where models are more adept at interpreting the cultural norms of their training data.
4.4 Analysis
Section 4.2 reveals a critical survey-based specialization pitfall: base models fine-tuned on sanitized survey datasets exhibit degraded performance even with context priming. This unexpected finding raises three questions that we expect to investigate:
  • (1) What role does social context actually play?
  • (2) Why do models struggle with contextual complexity?
  • (3) Is it more beneficial for a model to understand specific context or to imitate examples?
These questions emerge naturally from our specialization findings. The failure of fine-tuned models despite having full context suggests either that platform, topic and temporal context are less important than assumed, or that these models cannot effectively utilize those contextual information. This motivates our first investigation into whether social context is truly beneficial signal or noise. The stark performance gap between base and specialized models further suggests that certain contextual characteristics may be particularly challenging, leading to our second analysis of contextual complexity dimensions. Finally, the contrast between our context-rich approach and traditional few-shot learning methods—where specialized models were trained on examples rather than contextual understanding—compels us to directly compare these two information-providing strategies.
@@ 4.4 Analysis
Social Context is a Critical Signal. Our context ablation study (Table 4) confirms that social context is not noise but a vital signal for accurate opinion prediction. Removing all contextual metadata (No Context) results in the most substantial performance drop on average (-5.91%), followed closely by platform-specific information removal (-5.12%) and topical context removal (-4.52%). Temporal context shows the smallest impact, but its existence is important to provide backgrounds for questions need temporal information.
Notably, specialized fine-tuned models exhibit dramatically larger degradation across all ablation conditions. This severe degradation approaches the theoretical lower bound of uniform distribution performance, suggesting these models have become overly dependent on contextual signals. Since these models were fine-tuned on structured survey data in standardized formats, they struggle to adapt when deployed on natural social media polls when contextual information may be fragmented or entirely absent. While these models may excel in controlled survey environments, their heightened sensitivity to context removal in real-world social media settings reveals a limitation, potentially constraining their applicability across diverse online environments.
@@
Contextual Complexity. Our ablation study showed that removing social context severely degrades performance. To understand why this context is so critical, we investigate how its inherent complexity impacts model behavior. We define contextual complexity using three metrics: (1) character length of combined metadata; (2) language informality (measured by slang and emoji frequency); and (3) topic niche nature (inversely proportional to dataset percentage). As demonstrated in Figure 5, our findings reveal consistent negative correlations between model performance and all complexity metrics across evaluated systems. This indicates that as social contexts become longer, more informal, and more specialized, models systematically exhibit reduced prediction certainty. The analysis reveals distinct performance tiers, with closed-source models showing greater resilience, open-source models displaying moderate sensitivity, and specialization models exhibiting the most pronounced degradation under complexity. This pattern suggests that domain-specific fine-tuning on formal survey data creates brittleness when encountering the informal, less structured nature of social media polls. The hierarchical performance differences confirm that authentic social discourse’s rich, messy characteristics pose universal challenges for LLMs, with models fine-tuned on structured formats being particularly vulnerable to contextual complexity.
Table 4: Opinion distiribution prediction performance degradation from the full-context baseline. All scores represent the drop in 1-Wasserstein Distance (%). Abbreviations: Plat. (Platform), Topic. (topical), Temp. (Temporal), Ctx. (Context).
Model w/o Plat. w/o Topic. w/o Temp. No Ctx.
@@
o3-medium -3.03 -2.31 -1.02 -4.13
Gemini-2.5-Pro -2.37 -3.30 -0.98 -4.34
Claude-3.7-Sonnet -4.46 -3.91 -1.39 -5.53
GPT-4o -5.75 -3.39 -0.87 -4.92
GPT-4.1 -4.92 -3.36 -1.57 -5.30
@@
Deepseek-R1 -4.19 -3.55 -1.04 -5.98
Qwen2.5-32B -4.42 -3.81 -0.88 -6.24
Llama-3-70B -4.82 -4.78 -2.11 -5.19
Llama-4-17B -5.80 -5.33 -1.01 -6.08
Llama-2-13B -4.95 -5.26 -1.37 -5.92
Gemma-2-9B -5.37 -4.74 -1.26 -6.47
Mistral-7B -5.92 -4.84 -1.46 -6.55
@@
SubPop-Llama-3-70B -6.82 -6.68 -1.88 -6.95
SubPop-Llama-2-13B -6.94 -6.11 -2.10 -7.23
SubPop-Mistral-7B -6.89 -6.41 -2.96 -7.54
Average -5.12 -4.52 -1.46 -5.91
Contextual Priming vs. Few-Shot Learning. Our “survey-based specialization pitfall" finding (Section 4.2) raised a key question: do models benefit more from understanding a poll’s environment via contextual priming or from imitating examples via few-shot learning? To test this, we compared model performance in a zero-shot setting with full context against a zero-shot setting with no context and a few-shot setting with no context.
The results in Figure 6 are stark: contextual priming yields a substantial increase across all models, exceeding the impact of few-shot examples in most cases. Moreover, few-shot performance shows instability across models but contextual priming provides more consistent improvements. This finding suggests that the ability to situate a problem within its social context is a more critical capability than pattern-matching from isolated examples.
Figure 5: Performance degradation patterns across model types and complexity dimensions. The heatmap reveals a clear hierarchy: closed-source models maintain robustness while survey-specialized models exhibit severe brittleness to social media contextual complexity.
@@ 5 Error Analysis
We conduct an error analysis to categorize prediction reasoning errors  [^1] We consider the individual poll’s 1-Wass. $<$ 0.8 as error.. Using the LLM-as-a-judge, we identify three main categories of error type shown in Table 5.
Table 5: Analysis of Claude 3.7 Sonnet prediction errors categorized by failure modes: Platform Mis-Adaptation (42.5%) – misapplying knowledge about specific platforms; Cultural Misalignment (36.6%) – ignoring local and cultural contexts; Temporal Dislocation (20.0%) – mis-understanding about the temporal context.
@@
Temporal Dislocation Calling it "X" vs. "Twitter" (Reddit): Reasoning anchored to official rebrand timeline, underestimating persistent colloquial usage and real-world adoption resistance.
Platform Mis-Adaptation. A model can fail into platform mis-adaptation, where they misapplying their knowledge about certain platform, such as when analyzing a Reddit poll about paid TikTok subscriptions. It may apply a generic economic lens, predicting high tendency toward worth option based on the influencer economy. This reasoning ignores the context of Reddit, where the user tend to be negative of influencer monetization. The prediction is thus inverted from the ground truth which shows a much higher proportion of not worth it.
Cultural Misunderstanding. In the phone replacement case, models systematically misinterpret Chinese users’ pragmatic spending attitudes, failing to capture the genuine cost-conscious preferences common among Weibo users, which often contradict Western-centric consumer marketing narratives. While model recognizes about the technical enthusiasm in Weibo, their predictions fail to reflect the cost-conscious approach that characterizes technology adoption for many Chinese social media users.
Temporal Dislocation. This error happens when models fail to account for the lag between official changes and actual user adoption patterns. In the Twitter-to-X rebranding case, the model assumes that users would quickly adopt the new term. However, this prediction overlooks that Reddit users, in particular, demonstrated strong attachment to the familiar terminology, viewing continued use of the old name as both habitual behavior and subtle resistance. The model’s temporal reasoning thus misaligned with the gradual and reluctant pace of real-world language adoption.
@@ 6 Conclusion
We introduce MindVote, the first benchmark for public opinion prediction grounded in authentic social media discourse. Our comprehensive evaluation demonstrates the critical importance of assessing models in naturalistic, context-rich environments. We argue that the path to socially intelligent AI requires enhancing a model’s capacity for in-context reasoning. Our results demonstrate that models perform best when they can explicitly identify, weigh, and interpret the social cues present in the immediate context—a skill that requires flexible reasoning rather than memorized associations. MindVote provides the essential tool to guide and measure this necessary shift.

### v5

@@ 4.3 Performance Across Topics and Cultures
Performance varies by Topic. As shown in Table 3, model performance differs notably across topical domains. On average, models perform better on Social Issues and Lifestyle topics, which typically exhibit more structured discourse aligned with standard pre-training corpora. In contrast, performance declines in domains such as Technology and Entertainment, where specialized jargon, community-specific norms, and rapidly evolving vernacular are prevalent. This disparity indicates that models have difficulty adapting their reasoning from general knowledge to the nuanced linguistic and social norms of diverse online communities.
@@
Performance Reflects Cultural Origin. The analysis in Figure 4 reveals a Cultural Gap, showing models have a strong cultural alignment with their origin. Western models like Gemini-2.5-Pro excel on Reddit, while Chinese models such as DeepSeek-R1 dominate on Weibo. We confirmed this gap is primarily cultural, not linguistic, by evaluating translated content; the performance penalty from translation was minimal compared to the large drop across cultural sources. This systematic variation demonstrates a clear alignment effect, where models are more adept at interpreting the cultural norms of their training data.
4.4 Analysis
Section 4.2 highlights a pitfall of survey-based specialization: base models fine-tuned on sanitized survey datasets show degraded performance, even when context priming—adding relevant situational or background information to guide predictions—is applied (Doyle and Lee 2016). This unexpected finding raises three questions that we expect to investigate:
  • Does social context in social media serve as noise or a helpful signal?
  • How does increasing contextual complexity affect model performance, especially for survey-specialized models?
  • Does adding social contexts better enhance models’ opinion prediction compared with few-shot learning?
These questions emerge naturally from our specialization findings. The failure of fine-tuned models, even when provided with full context, suggests two possibilities: either (1) platform and topic are less important than previously assumed, or (2) these models are unable to effectively utilize such contextual information. This motivates our first investigation into whether social context is truly beneficial signal or noise. Moreover, the suboptimal performance of specialized models reflects their difficulty in processing social context, motivating our second analysis on factors contributing to context complexity. Third, because our context-rich approach relies on contextual understanding rather than example-based training as in few-shot learning, a direct comparison is needed to assess the effectiveness of both strategies.
@@ 4.4 Analysis
Social Context is a Critical Signal. Our context ablation study (Table 4) confirms that social context is not noise but a vital signal for accurate opinion prediction. Removing all contextual metadata (No Context) results in the most substantial performance drop on average (-5.91%), followed closely by platform-specific information removal (-5.12%) and topical context removal (-4.52%).
Notably, specialized fine-tuned models exhibit dramatically larger degradation across all ablation conditions. This severe degradation approaches the theoretical lower bound of uniform distribution performance, suggesting these models have become overly dependent on contextual signals used for priming. Since these models were fine-tuned on structured survey data in standardized formats, they struggle to adapt when deployed on natural social media polls when contextual information may be fragmented or entirely absent. While these models may excel in controlled survey environments, their heightened sensitivity to context removal in real-world settings reveals a limitation, constraining their applicability across diverse online environments.
Model w/o Plat. w/o Topic. No Ctx.
@@
o3-medium -3.03 -2.31 -4.13
Gemini-2.5-Pro -2.37 -3.30 -4.34
Claude-3.7-Sonnet -4.46 -3.91 -5.53
GPT-4o -5.75 -3.39 -4.92
GPT-4.1 -4.92 -3.36 -5.30
@@
Deepseek-R1 -4.19 -3.55 -5.98
Qwen2.5-32B -4.42 -3.81 -6.24
Llama-3-70B -4.82 -4.78 -5.19
Llama-4-17B -5.80 -5.33 -6.08
Llama-2-13B -4.95 -5.26 -5.92
Gemma-2-9B -5.37 -4.74 -6.47
Mistral-7B -5.92 -4.84 -6.55
@@
SubPop-Llama-3-70B -6.82 -6.68 -6.95
SubPop-Llama-2-13B -6.94 -6.11 -7.23
SubPop-Mistral-7B -6.89 -6.41 -7.54
Average -5.12 -4.52 -5.91
Table 4: Opinion distribution prediction performance degradation from the full-context baseline. All scores represent the drop in 1-Wasserstein Distance (%). Abbreviations: Plat. (Platform), Topic. (topical), Ctx. (Context).
Contextual Complexity. While our ablation study confirms that social context is critical for performance, its inherent complexity presents a significant challenge. To dissect this, we investigate how model performance degrades as complexity increases about our annotated context. We define contextual complexity using three metrics: (1) context length by tokens of combined metadata; (2) language informality in the context; and (3) niche topic which constitutes a small fraction of the overall dataset. Our analysis in Figure 5 reveals that increasing complexity in any of these areas has negative correlations with the 1-Wass. performance.
This effect is most pronounced for our specialization models. Specifically, these models, fine-tuned on structured and formal survey data, exhibit strong performance degradation when faced with longer contexts, higher degrees of informality, and more niche topics. This demonstrates a critical brittleness: while trained to be domain experts, their reliance on formal data structures makes them particularly vulnerable to the unstructured, multifaceted, and messy nature of authentic social media discourse.
Figure 5: Correlation patterns between model’s 1-Wass. performance and complexity dimensions. Survey-specialized models exhibit strong brittleness to social media contextual complexity.
Contextual Priming vs. Few-Shot Learning. Our “survey-based specialization pitfall” finding (Section 4.2) raised a key question: do models benefit more from understanding a poll’s environment via contextual priming or from imitating examples via few-shot learning? To evaluate this, we compared the model’s performance across three distinct settings. The first was a zero-shot setting with context priming, which followed our default context annotation pipeline. We contrasted this against two settings without context priming: a standard zero-shot setting without context priming, same as we have done in Table 4 and a few-shot (1-4 examples) setting. For the few-shot configuration, we supplied the model with examples predicted from Claude-3.7, where each example contained a poll, its distribution prediction, and the corresponding reasoning steps.
The results in Figure 6 are stark: contextual priming delivers a substantial performance boost across all models, often exceeding the gains achieved through few-shot learning. In contrast, the benefit from few-shot examples is unreliable. As the figure shows, performance does not reliably increase with the number of examples; for instance, the average 4-shot performance of open-sourced models is worse than that of 3-shot performance. This discrepancy suggests that the ability to situate a problem within its broader social context is a more critical and robust capability than simple pattern-matching from isolated examples.
@@ 5 Error Analysis
We conduct an error analysis to categorize prediction reasoning errors  [^1] We consider the individual poll’s 1-Wass. << 0.8 as error.. Using the LLM-as-a-judge, we identify three main categories of error type shown in Table 5.
@@
Temporal Dislocation Calling it ”X” vs. ”Twitter” (Reddit): Reasoning anchored to official rebrand timeline, underestimating persistent colloquial usage and real-world adoption resistance.
Table 5: Analysis of Claude 3.7 Sonnet prediction errors categorized by failure modes: Platform Mis-Adaptation (42.5%) – misapplying knowledge about specific platforms; Cultural Misalignment (36.6%) – ignoring local and cultural contexts; Temporal Dislocation (20.0%) – mis-understanding about the temporal information.
Platform Mis-Adaptation. In the TikTok subscriptions case, the model applies a generic economic lens, predicting high tendency toward worth option based on the influencer economy. This reasoning ignores the context of Reddit, where the user tend to be negative of influencer monetization. The prediction is thus inverted from the ground truth which shows a much higher proportion of not worth it.
Cultural Misunderstanding. In the phone update case, models systematically misinterpret Chinese users’ pragmatic spending attitudes, failing to capture the genuine cost-conscious preferences common among Weibo users, which often contradict Western-centric consumer marketing narratives. While model recognizes about the technical enthusiasm in Weibo, their predictions fail to reflect the cost-conscious approach that characterizes technology adoption for many Chinese social media users.
Temporal Dislocation. In the Twitter-to-X rebranding case, the model assumes that users would quickly adopt the new term. However, this prediction overlooks that Reddit users, in particular, demonstrated strong attachment to the familiar terminology, viewing continued use of the old name as both habitual behavior and subtle resistance. The model’s temporal reasoning thus misaligned with the gradual and reluctant pace of real-world language adoption.
@@ 6 Conclusion
We introduce MindVote, the first benchmark for public opinion distribution prediction in social media. Our comprehensive evaluation demonstrates the critical importance of assessing models in naturalistic, context-rich environments. We argue that the path to socially intelligent AI requires enhancing a model’s capacity for in-context reasoning. Our results demonstrate that models perform best when they can explicitly identify, weigh, and interpret the social cues present in the immediate context—a skill that requires flexible reasoning rather than memorized associations. MindVote provides the essential tool to guide and measure this necessary shift.
7 Limitation
While MindVote represents a significant advance in ecological validity, our benchmark is limited to two platforms, Reddit and Weibo, which, despite their scale and cultural diversity, do not capture the full spectrum of social media ecosystems. Platforms such as Twitter/X, TikTok, Instagram, and region-specific networks exhibit distinct interaction patterns, demographic compositions, and content modalities that may influence opinion formation differently. Future work should expand to additional platforms to ensure comprehensive evaluation of LLMs’ social reasoning capabilities across diverse digital environments.